OpenStudy (anonymous):

suppose z=(y^4)(x^3). at the instant when x=1 and y=3, x is decreasing at 2 units/sec and y is increasing at 2 units/sec. how fast is z changing at this instant? is z increasing or decreasing?

OpenStudy (anonymous):

Based on the chain rule:\[ \frac{dz}{dt}=\frac{\partial z}{\partial x}\frac{dx}{dt}+\frac{\partial z}{\partial y}\frac{dy}{dt} \]

OpenStudy (anonymous):

"x is decreasing at 2 units/sec"\[ \frac{dx}{dt}=-2 \]"y is increasing at 2 units/sec"\[ \frac{dy}{dt}=2 \]

OpenStudy (anonymous):

@chels.408 Do you get it?

OpenStudy (anonymous):

not really..

OpenStudy (anonymous):

Do you know partial derivatives?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

You must know multivariable calculus to do this problem, so you need to know partial derivatives.

OpenStudy (anonymous):

\[ z=(y^4)(x^3) \]You can do partial derivative with respect to \(x\)?

OpenStudy (anonymous):

\[ z=f(x,y)=(y^4)(x^3) \]\[ \frac{\partial z}{\partial x}=f_x(x,y)=3(y^4)(x^2) \]

OpenStudy (anonymous):

I just treat \(y\) as a constant and differentiate with respect to \(x\). Do you understand?

OpenStudy (anonymous):

oooh I see. thanks

OpenStudy (anonymous):

So find \(f_x(1,3)\) and \(f_y(1,3)\)

OpenStudy (anonymous):

\[ z'=f_x(1,3)x'+f_y(1,3)y' \] This should make it a bit more clear.

OpenStudy (anonymous):

"x is decreasing at 2 units/sec"\[ x'=-2 \]"y is increasing at 2 units/sec"\[ y'=2 \]

OpenStudy (anonymous):

@chels.408 Can you do it now?

OpenStudy (anonymous):

yes thank you